# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v1

@@ Appendix A System Setup Details
The H&R system is equipped with state-of-the-art technology specifically chosen to facilitate seamless and precise interactions between human operators and robotic mechanisms, including a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras as shown in Figure 9(a) 9(b).
@@ B.1 Coordinate System Alignment
(a) Different positions. We positioned the block at two distinct locations in the scene (top-left and top-right of the image) and observed that the model was able to complete the tasks successfully in both cases. Additionally, we placed the box at two different positions on the table, and the model was able to complete the tasks effortlessly.
(b) Unseen object and trajectory. We first tested the model with unseen objects, and the object in the human demonstration video differed from the actual object (top-left image). The results showed that the model successfully completed the task without any noticeable impact. Furthermore, we evaluated the model on grasping tasks involving unseen objects (top-right and bottom-left images). The results demonstrated that the model could successfully complete tasks with objects it had not encountered before. Lastly, we tested the model on unseen tasks with unseen objects (bottom-right image). The results revealed that our model exhibited strong generalization capabilities and successfully completed the task.
Figure 10: Qualitative visualization of generated results and robot actions
@@ C.1 Task Details
Our H&R dataset contains 6 base tasks and 6 long-horizon tasks. The definition of tasks in dataset is listed below.
@@
  • Pinch a cup: The cup can initially be placed anywhere on the table. The goal of this task is to pick up the cup and place it in another location.
  • Pinch a cube: The cube comes in two colors (red and green), and it can initially be placed within a designated area. The goal of this task is to pick up the cube and place it on a plate.
  • Pinch a pencil: The pencil comes in two colors (red and black), the goal is to pick up the pencil and place it on a plate.
  • Push a box: The box comes in two colors (red and green), and its initial position of it is random. The goal is to push the box from left to right
  • Pull a plate: The goal is to pull the plate from bottom to top.
  • Push a plate: The goal is to push the plate from top to bottom.
@@
  • Pinch both cubes: There are two red cubes on the table, and the goal is to pick up each cube individually and place them onto the plate.
  • Pinch the designated cube: There is a red cube and a green cube on the table, and the goal is to pick up one of the cubes onto the plate.
  • Pinch to the designated plate: There is a red cube and two plates on the table, and the goal is to pick up the cubes onto one of the plates.
  • Pinch the cube and pull the plate: The combination of pinch a cube and pull the plate. The goal is to first pick the cube onto the plate and pull the plate.
  • Pull the plate and Pinch the cube: The combination of pull the plate and pinch a cube. After pulling the plate, the cube should be placed in the plate’s new position.
@@ C.2 Episode Details
H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X [43] standards:
@@
Appendix D Baseline Details
Here we talk about how we add conditions to ACT [71] in detail. We input the observation information of the robotic arm into the encoder. Then, we add the features of the next $n$ frames of the human hand as a condition to the position embedding and input them into the decoder. Specifically, similar to the method used for processing the original robot images, we input human images into ResNet18 [22] and obtain a feature map of size 300x512. We then pass it through a shared MLP to obtain a feature of size 512, which is added to the position embedding and fed into the Transformer decoder. Both the ResNet18 [22] and MLP are trainable.

### v2

@@ Appendix A System Setup Details
The H&R system is equipped with state-of-the-art technology specifically chosen to facilitate seamless and precise interactions between human operators and robotic mechanisms, including a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras as shown in Figure 7(a) 7(b).
@@ B.1 Coordinate System Alignment
B.2 System stability
Accuracy. There are two main factors affecting control accuracy: one is the precision of the data source, i.e., the accuracy of hand gesture recognition, and the other is the accuracy loss caused by mapping hand movements to the robotic arm. Our gesture recognition data comes from the Meta Quest 3, and we found that under sufficient lighting conditions, it provides highly accurate hand tracking.
A traditional mapping approach is to map the wrist position to the end-effector position of robot. However, this creates a topological inconsistency, leading to accuracy errors. A more topologically consistent approach is to map the midpoint between the thumb and index fingertip. While this method preserves structural consistency, it introduces jitter due to the differing speeds of the thumb and index finger when grasping objects. Thus, we track the Metacarpophalangeal Joint (MCP) of the index finger, ensuring both the prevention of jitter and the maintenance of accuracy.
Jitter Handling. We set the control mode to serial, making the system’s sensitivity depend on execution speed. By allowing 100–300ms latency, we trade speed for stability, reducing jitter and ensuring smoother motion.
Figure 9: Visualization of generalization evaluation process. To accelerate inference, we skip the video generation process when performing the handwriting task.
@@ C.1 Task Details
Our H&R dataset contains 8 base tasks and 6 long-horizon tasks. The definition of our tasks in our dataset is listed below.
@@
  • Pick and place the cup: The cup can initially be placed anywhere on the table. The goal of this task is to pick up the cup and place it in another location.
  • Pick and place the cube: The cube comes in two colors (red and green), and it can initially be placed within a designated area. The goal of this task is to pick up the cube and place it on a plate.
  • Pick and place pencil: The pencil comes in two colors (red and black), the goal is to pick up the pencil and place it on a plate.
  • Push the box: The box comes in two colors (red and green), and its initial position of it is random. The goal is to push the box from left to right
  • Pull the plate: The goal is to pull the plate from bottom to top.
  • Push the plate: The goal is to push the plate from top to bottom.
  • Pick up and hold the brush upright: The goal is to pick up the brush and upright it.
  • Handwriting: Play data, write aimlessly on a blank desk.
@@
  • Pick both cubes: There are two red cubes on the table, and the goal is to pick up each cube individually and place them onto the plate.
  • Pick the designated cube: There is a red cube and a green cube on the table, and the goal is to pick up one of the cubes onto the plate.
  • Pick to the designated plate: There is a red cube and two plates on the table, and the goal is to pick up the cubes onto one of the plates.
  • Pick the cube and pull the plate: The combination of pinch a cube and pull the plate. The goal is to first pick the cube onto the plate and pull the plate.
  • Pull the plate and pick the cube: The combination of pull the plate and pinch a cube. After pulling the plate, the cube should be placed in the plate’s new position.
@@ C.2 Episode Details
H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X standards:
@@
Appendix D Experimental Details
D.1 Experimental Details
D.2 Baseline Details
RVCD Policy. Here we talk about how we add conditions to ACT [72] in detail. We input the observation information of the robotic arm into the encoder. Then, we add the features of the next $n$ frames of the human hand as a condition to the position embedding and input them into the decoder. Specifically, similar to the method used for processing the original robot images, we input human images into ResNet18 [22] and obtain a feature map of size 300x512. We then pass it through a shared MLP to obtain a feature of size 512, which is added to the position embedding and fed into the Transformer decoder. Both the ResNet18 [22] and MLP are trainable.
XSkill. We replicated xSkill in a real-world setting, where the authors did not fully open-source the code. As a result, we completed the remaining code based on the code from the simulation environment. During the skill discovery and labeled dataset phases, we used the default settings, except for adjusting the num frames parameter of the frame sampler from 100 to 50 to reduce GPU memory consumption. We train this stage for 300 epochs.
In the skill transfer composing phase, we found that performance of XSkill was quite poor. Although the behavior cloning loss decreased significantly, the diffusion policy barely demonstrated any task execution ability after denoising. Since the XSkill [69] paper mentions that any imitation learning policy can be used with the XSkill framework, we replaced the diffusion policy with an action head the same as Human2Robot, a simple MLP, for training. During training, we adjusted the prediction horizon to 32, resized the pretraining to 160x120 to maintain consistency with the previous training process, and set the vision feature dimension to 64, in line with the simulation environment. We trained this stage for a total of 2000 epochs until the loss showed no significant changes.